OpenStudy (anonymous):

How would I cuberoot 16x^5y^9?

OpenStudy (anonymous):

2xy^3cubert(8x^2)

OpenStudy (anonymous):

\[\sqrt[3]{16x ^{5}y ^{9}}\] =\[\sqrt[3]{16}x^{5/3}y ^{9/3}\] =\[2\sqrt[3]{2}x ^{5/3}y ^{3}\]